# Self-Improving Skill Learning for Robust Skill-based Meta-Reinforcement Learning

### v3

@@ D.3 Random Noise Injection
While Gaussian noise is a widely used approach for degrading demonstration quality in offline RL studies, it does not fully capture the diverse and unstructured nature of real-world noise such as sensor failures, occlusions, or actuator malfunctions. In our main experiments, we adopted multi-level Gaussian noise for two reasons: (1) it is a standard and accepted method for systematically degrading demonstration quality, and (2) it enables controlled analysis of robustness under varying degrees of skill degradation. Notably, in domains such as Kitchen and Office, sufficiently high levels of Gaussian noise render learned skills nearly unusable, effectively mimicking real-world failure scenarios in precision control tasks and providing a highly challenging regime for evaluating robustness.
@@ D.4 Diverse Sub-Optimal Dataset
Beyond injecting noise into expert demonstrations, we also considered datasets of diverse quality sub-optimal demonstration. Our decision to focus on noisy expert trajectories stems from the specific challenge we aim to address. Unlike offline-RL, which typically assumes access to reward-labeled trajectories and aims to improve performance from sub-optimal data, our setup considers reward-free offline data where noise degrades originally near-optimal demonstrations. This setting better reflects our goal of studying how exploration can help enhance skill learning when clean supervision is unavailable. To test whether robustness hold across different data qualities, we conducted additional experiments in the Kitchen domain using three dataset types: expert (return == 4), medium (return $\approx$ 2), and random (collected from a uniformly random policy). As summarized in Table D.3, SISL consistently outperforms other methods across all dataset types, further validating its robustness to data sub-optimality.
@@ Appendix E Computational Complexity
Despite SISL’s significant performance improvements on noisy demonstrations through self-improving skill refinement, it is important to consider the computational trade-offs. Simultaneously training both low-level and high-level policies in SISL can increase computation time. During meta-train, SISL requires approximately 16% more time per iteration (SiMPL: 15.1s vs SISL: 17.6s in Kitchen). However, as shown in Fig. 6, SiMPL does not improve under high-noise settings even with extended training, whereas SISL continues to improve, justifying this computational overhead. Importantly, considering that the main goal of meta-RL is adaptation to unseen tasks, SISL does not require additional computational cost during the fine-tuning phase and maintains similar computation time as SiMPL in the meta-test phase while still achieving superior performance compared to baseline algorithms.
@@ Appendix F Further Analysis Results for the SISL Framework
In this section, we provide a more detailed analysis and visualization results of the skill refinement process in the SISL. The analysis performed includes the evolution of the mixing coefficient $\beta$ in Section F.1, visualization of the refined skills and the corresponding skill sequence visualization in Section F.2, task representation in Section F.3, and the learning stability of the reward model during the meta-train phase in Section F.4.
@@ F.1 Evolution of The Mixing Coefficient $\beta$
Fig. F.1 illustrates the evolution of the mixing coefficient $\beta$ during the meta-train phase across all environments and noise levels. Initially, low $\beta$ values reflect reliance on offline datasets for skill learning, particularly in high-return environments like Kitchen and Office, where training starts with $\beta$ values close to zero. This approach prevents performance degradation by avoiding early dependence on lower-quality online samples, while ensuring gradual and stable changes in the skill distribution. As training progresses, the quality of online samples improves, leading to a gradual increase in $\beta$, which facilitates greater utilization of online data for skill refinement. For offline datasets with higher noise levels, $\beta$ converges to higher values. Consequently, SISL learns increasingly effective skills as training proceeds, achieving superior performance on unseen tasks, underscoring the importance of SISL’s ability to dynamically balance the use of offline and online data based on dataset quality.
@@ F.4 Stability of Reward Model Learning
Unlike offline-RL, which typically relies on reward-labeled datasets, skill-based RL operates on reward-free offline data. In SISL, only the skills are learned from the noisy offline trajectories, while the reward model and the high-level policy are trained online using clean, noise-free training tasks, as in SiMPL. As shown in Eq. (4), the reward model is trained using transitions sampled from $\mathcal{B}^{i}_{\text{imp}}$ and $\mathcal{B}^{i}_{\text{on}}$, which contain only online trajectories collected by interacting with training task $i$. Since these buffers contain no offline data, the reward model is unaffected by offline noise. To further support this point, we present Table F.1, which shows that the MSE of the reward model remains consistently low across different offline noise settings in the Kitchen environment. These results indicate that the reward model maintains high accuracy and continues to support effective relabeling even when offline skill pretraining is noisy.
@@ Appendix G Additional Ablation Studies
In this section, we conduct additional ablation studies for Kitchen and Maze2D environments across all noise levels. These studies include component evaluation and SISL’s skill refinement-related hyperparameters discussed in Section G.1, prioritization temperature $T$ in Section G.2, the KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ for the skill-improvement policy in Section G.3, and additional component evaluation on RND and re-initialization in Section G.4.
@@ G.4 Additional Component Evaluation
The SISL w/o Re-Initialize refers to a variant that continues training the high-level policy $\pi_{h}$ without re-initializing it at each $K_{\text{iter}}$ during the meta-training process. In SISL, the skill model is periodically updated every $K_{\text{iter}}$ steps, which effectively changes the environment dynamics observed by the high-level policy. Continuing to train the same high-level policy across different skill sets introduces non-stationarity, leading to instability. To address this, we reset both the policy parameters and the buffer at each skill update so that the high-level policy can re-learn from scratch under a new, stable MDP defined by the updated skills. This technique is consistent with practices in continual and safe reinforcement learning, where re-initialization is often used to manage sudden changes in task dynamics (Kim et al., 2023; Kong et al., 2024). To empirically validate this decision, we conducted ablation experiments on high-level policy re-initialization. The results show that removing re-initialization leads to a substantial performance drop, supporting the necessity of this design choice.

### v4

@@ D.3 Random Noise Injection
While Gaussian noise is a widely used approach for degrading demonstration quality in offline RL studies, it does not fully capture the diverse and unstructured nature of real-world noise such as sensor failures, occlusions, or actuator malfunctions. In our main experiments, we adopted multi-level Gaussian noise for two reasons: (1) it is a standard and accepted method for systematically degrading demonstration quality, and (2) it enables controlled analysis of robustness under varying degrees of skill degradation. Notably, in domains such as Kitchen and Office, sufficiently high levels of Gaussian noise render learned skills nearly unusable, mimicking real-world failure scenarios in precision control tasks and providing a highly challenging regime for evaluating robustness.
@@ D.4 Diverse Sub-Optimal Dataset
Beyond injecting noise into expert demonstrations, we also considered datasets of diverse quality sub-optimal demonstrations. Our decision to focus on noisy expert trajectories stems from the specific challenge we aim to address. Unlike offline-RL, which typically assumes access to reward-labeled trajectories and aims to improve performance from sub-optimal data, our setup considers reward-free offline data where noise degrades originally near-optimal demonstrations. This setting better reflects our goal of studying how exploration can help enhance skill learning when clean supervision is unavailable. To test whether robustness holds across different data qualities, we conducted additional experiments in the Kitchen domain using three dataset types: expert (return == 4), medium (return $\approx$ 2), and random (collected from a uniformly random policy). As summarized in Table D.3, SISL consistently outperforms other methods across all dataset types, further validating its robustness to data sub-optimality.
@@ Appendix E Computational Complexity
Table E.1 summarizes SISL’s computational complexity, showing that the meta-train time increases by about 16% relative to SiMPL, which we view as a modest overhead given the performance gains. For a fair comparison, we keep the amount of training and interaction identical to SiMPL as described in Section 5.1, so the number of samples and policy updates does not exceed that of SiMPL.
Table E.2 summarizes the complexity of SISL ablations. “SISL w/o $P_{\mathcal{B}_{\text{off}}}$” and “SISL w/o $\mathcal{B}_{\text{off}}$” do not perform reward model training and only add skill refinement compared to SiMPL, increasing computation time by about 10%. In contrast, “SISL w/o $\pi_{\text{imp}}$”, “SISL w/o $\mathcal{B}_{\text{on}}$”, and full SISL additionally perform reward model training, incurring an additional increase of about 3% in computation time. In particular, “SISL w/o $\pi_{\text{imp}}$” differs from full SISL by only 2.3% in training time, indicating that introducing $\pi_{\text{imp}}$ does not lead to a substantial increase in computation cost.
Thus, the 16% overhead mainly arises from skill refinement and reward model training rather than additional policy learning. This overhead is modest in light of the performance gains: as shown in Fig. 6, SiMPL does not improve under high-noise settings even with extended training, whereas SISL continues to improve, justifying this computational overhead. Importantly, considering that the main goal of meta-RL is adaptation to unseen tasks, SISL does not require additional computational cost during the fine-tuning phase and maintains similar computation time as SiMPL in the meta-test phase while still achieving superior performance compared to baseline algorithms.
Table E.1: Comparison of total training time for SiMPL and SISL in Kitchen(Expert).
Algorithm Training time (h) $\Delta$ time vs. SiMPL (%)
SiMPL 42.16 -
SISL 48.96 +16.1
Table E.2: Comparison of total training time for SISL ablation variants in Kitchen(Expert).
Algorithm Training time (h) $\Delta$ time vs. SiMPL (%)
SISL w/o $\mathcal{B}_{\text{off}}$ 46.23 +9.6
SISL w/o $P_{\mathcal{B}_{\text{off}}}$ 46.71 +10.7
SISL w/o $\pi_{\text{imp}}$ 47.98 +13.8
SISL w/o $\mathcal{B}_{\text{on}}$ 48.81 +15.7
@@ Appendix F Further Analysis Results for the SISL Framework
In this section, we provide a more detailed analysis and visualization results of the skill refinement process in the SISL. It includes the evolution of the mixing coefficient $\beta$ in Section F.1, visualization of the refined skills and the corresponding skill sequence visualization in Section F.2, task representation in Section F.3, learning stability of the reward model during the meta-train phase in Section F.4, and comparison of skill trajectory with baseline algorithms in Section F.5.
@@ F.1 Evolution of The Mixing Coefficient $\beta$
Fig. F.1 illustrates the evolution of the mixing coefficient $\beta$ during the meta-train phase across all environments and noise levels. Initially, low $\beta$ values reflect reliance on offline datasets for skill learning, particularly in high-return environments like Kitchen and Office, where training starts with $\beta$ values close to zero. This approach prevents performance degradation by avoiding early dependence on lower-quality online samples, while ensuring gradual and stable changes in the skill distribution. As training progresses, the quality of online samples improves, leading to a gradual increase in $\beta$, which leverages online data for skill refinement. For offline datasets with higher noise levels, $\beta$ converges to higher values. Consequently, SISL learns increasingly effective skills as training proceeds, achieving superior performance on unseen tasks, underscoring the importance of SISL’s ability to dynamically balance the use of offline and online data based on dataset quality.
@@ F.4 Stability of Reward Model Learning
Unlike offline-RL, which typically relies on reward-labeled datasets, skill-based RL operates on reward-free offline data. In SISL, only the skills are learned from the noisy offline trajectories, while the reward model and the high-level policy are trained online using clean, noise-free training tasks, as in SiMPL. As shown in Eq. (4), the reward model is trained using transitions sampled from $\mathcal{B}^{i}_{\text{imp}}$ and $\mathcal{B}^{i}_{\text{on}}$, which contain only online trajectories collected by interacting with training task $i$. Since these buffers contain no offline data, the reward model is unaffected by offline noise. To further support this point, we present Table F.1, which shows the MSE of the reward model remains consistently low across different offline noise settings in the Kitchen environment. These results indicate that the reward model maintains high accuracy and continues to support effective relabeling even when offline skill pretraining is noisy.
@@
F.5 Comparison of Skill Trajectory Visualizations
To visually demonstrate the skill difference between the baseline algorithm and SISL, we have added a visualization comparison of the skill trajectories for SPiRL, SiMPL, and SISL. Specifically, Fig. F.5 and F.6 present the skill trajectory visualizations of the microwave-opening and bottom-burner control subtasks in the Kitchen($\sigma=0.3$). The SPiRL and SiMPL fail to complete these subtasks due to noise in the learned skills, either do not succeed in opening the microwave door or fail to properly reach the bottom-burner switch. In contrast, SISL successfully grasps and opens the microwave door and correctly manipulates the bottom-burner. These results highlight the substantial impact of noise in offline demonstrations on subtask performance, and demonstrate that SISL progressively refines skills even in noisy environments, successfully solving most given subtasks.
Figure F.5: Comparison of skill trajectory visualizations for the microwave-opening subtask in Kitchen($\sigma=0.3$) across algorithms.
Figure F.6: Comparison of skill trajectory visualizations for the bottom-burner control subtask in Kitchen($\sigma=0.3$) across algorithms.
@@ Appendix G Additional Ablation Studies
In this section, we conduct additional ablation studies for Kitchen and Maze2D environments across all noise levels. These studies include component evaluation and SISL’s skill refinement-related hyperparameters discussed in Section G.1, prioritization temperature $T$ in Section G.2, the KLD coefficient $\lambda^{\text{kld}}_{\text{imp}}$ for the skill-improvement policy in Section G.3, additional component evaluation on RND and re-initialization in Section G.4, skill refinement interval $K_{\text{iter}}$ in Section G.5, and comparison with Goal-Conditioned RL in Section G.6.
@@ G.4 Additional Component Evaluation
The SISL w/o Re-Initialize refers to a variant that continues training the high-level policy $\pi_{h}$ without re-initializing it at each $K_{\text{iter}}$ during the meta-training process. In SISL, the skill model is periodically updated every $K_{\text{iter}}$ steps, which effectively changes the environment dynamics observed by the high-level policy. Continuing to train the same high-level policy across different skill sets introduces non-stationarity, leading to instability. To address this, we reset both the policy parameters and the buffer at each skill update so that the high-level policy can re-learn from scratch under a new, stable MDP defined by the updated skills. This technique is consistent with practices in continual and safe reinforcement learning, where re-initialization is often used to manage sudden changes in task dynamics (Kim et al., 2023; Kong et al., 2024). To empirically validate this decision, we conducted ablation experiments on high-level policy re-initialization. The results show that removing re-initialization degrades performance, supporting the necessity of this design choice.
@@
G.5 Skill Refinement Interval $K_{\textnormal{iter}}$
As shown in Fig. 6, the high-level policy $\pi_{h}$ in the Kitchen environment typically requires at least 1K iterations within each interval to improve and converge, implying that if $K_{\text{iter}}$ is too small, skill refinement is triggered before $\pi_{h}$ has sufficiently adapted to the current skill library. Based on this observation, we used $K_{\text{iter}}=2000$ as the default setting in the Kitchen environment. To validate this intuition more rigorously, we conducted hyperparameter search over $K_{\text{iter}}\in\{100,500,1000,2000,5000\}$ in the Kitchen ($\sigma=0.3$) setting. The results confirm our hypothesis: with $K_{\text{iter}}=100$, $\pi_{h}$ is updated before it can meaningfully exploit the skills, leading to a substantial performance drop. Conversely, when $K_{\text{iter}}=5000$, $\pi_{h}$ converges well but skill refinement happens too infrequently, reducing the overall performance gain. These findings support the reasoning behind our chosen interval.
Figure G.4: Impact of the skill refinement interval $K_{\text{iter}}$ in Kitchen($\sigma=0.3$)
G.6 Comparison with Goal-Conditioned RL
We first note that standard goal-conditioned RL (GCRL) assumes explicit, environment-provided goals and single-task learning, whereas in our meta-RL setting no goals are given and the agent must infer task while learning a policy that generalizes across many tasks. Despite these differences, to provide an empirical reference point, we modify our environments to explicitly provide goals and evaluate a representative GCRL method, HER (Andrychowicz et al., 2017) with SAC, trained directly on the test tasks without offline data. As presented in Table G.2, HER achieves notably lower performance than SISL in Maze2D and fails to make progress in AntMaze, where the success rate remains at zero. Even with explicit goals, these sparse-reward long-horizon tasks appear difficult to solve without pretrained skills, which helps contextualize the advantages of our framework.
Table G.2: Final performance of goal-conditioned RL (HER) given a target goal.
Environment SAC+HER SISL
[…]
Maze2D($\sigma=0.5$) $\mathbf{0.89}$$\pm$0.03
Maze2D($\sigma=1.0$) $\mathbf{0.93}$$\pm$0.05
Maze2D($\sigma=1.5$) $\mathbf{0.99}$$\pm$0.02
AntMaze(Expert) $\pm0.00$ $\mathbf{0.81}$$\pm$0.08
AntMaze($\sigma=0.5$) $\mathbf{0.82}$$\pm$0.05
AntMaze($\sigma=1.0$) $\mathbf{0.60}$$\pm$0.02
AntMaze($\sigma=1.5$) $\mathbf{0.41}$$\pm$0.01